OpenStudy (anonymous):

1.03g of H2O to moles?

OpenStudy (anonymous):

And 0.07 to moles

OpenStudy (chmvijay):

what is molecular weight of water can you tell

OpenStudy (anonymous):

No im not sure how to do it

OpenStudy (chmvijay):

Molar mass or weight of H2O = 2* atomic mass of H+Atomic mass of O

OpenStudy (chmvijay):

hello say something

OpenStudy (anonymous):

Is it 18.02?

OpenStudy (chmvijay):

yaaa ur right molecular weight of any molecule = 1 mole of that molecule that is 18.02 gram of H2O = 1 mole then 1.03 gram of H2O =1*1.03 / 18.02 can u find now

OpenStudy (chmvijay):

did you get that ????

OpenStudy (anonymous):

I'm a little confused on how to find that part

OpenStudy (chmvijay):

which part

OpenStudy (anonymous):

Well how do i set it up to find the answer in moles

OpenStudy (chmvijay):

1 mole of any substance is equal to its molecular weight right

OpenStudy (chmvijay):

just do cross multiplication u get it 18.02 gram of H2O = 1 mole then 1.03 gram of H2O = X mole find X

OpenStudy (anonymous):

I got 17.495

OpenStudy (chmvijay):

nooo its not correct

OpenStudy (chmvijay):

X= 1* 1.03 / 18.02 = ??

OpenStudy (anonymous):

0.057?

OpenStudy (chmvijay):

yaaa right

OpenStudy (chmvijay):

 remeber always this